# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ Abstract.
Recent gain in popularity of AI conversational agents has led to their increased use for improving productivity and supporting well-being. While previous research has aimed to understand the risks associated with interactions with AI conversational agents, these studies often fall short in capturing the lived experiences. Additionally, psychological risks have often been presented as a sub-category within broader AI-related risks in past taxonomy works, leading to under-representation of the impact of psychological risks of AI use. To address these challenges, our work presents a novel risk taxonomy focusing on psychological risks of using AI gathered through lived experience of individuals. We employed a mixed-method approach, involving a comprehensive survey with 283 individuals with lived mental health experience and workshops involving lived experience experts to develop a psychological risk taxonomy. Our taxonomy features 19 AI behaviors, 21 negative psychological impacts, and 15 contexts related to individuals. Additionally, we propose a novel multi-path vignette based framework for understanding the complex interplay between AI behaviors, psychological impacts, and individual user contexts. Finally, based on the feedback obtained from the workshop sessions, we present design recommendations for developing safer and more robust AI agents. Our work offers an in-depth understanding of the psychological risks associated with AI conversational agents and provides actionable recommendations for policymakers, researchers, and developers.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Figure 1. Overview of our two-phase study. In the first phase, we conducted a survey study (N=283), using different analysis methods we created the Risk Taxonomy, comprising of Behavior, Impact, Context and their interplay. From these insights, we designed multi-path vignette and used it as a tool for the focus groups/workshops. In the second phase, we developed the future AI design recommendations through the workshop sessions with lived experience experts (N=7, Sessions=3).
@@
Since the late $\mathrm{20^{th}}$ century, advancements in technology, including personal computers and social media platforms, have enabled individuals to enhance productivity, support their well-being, and connect with other individuals (Dedrick et al., 2013; Sims et al., 2017; Hall et al., 2021; Shensa et al., 2016; Naslund et al., 2016). More recently, generative AI tools such as ChatGPT have gained in popularity that has led to an unprecedented growth in the number of individuals using these platforms to support their various productivity and well-being needs (Lebow, 2023; Afshar, 2024). Due to the natural conversational interface built on top of the underlying generative AI models that resemble human conversations, recent reports indicate that the market conversational AI agents is projected to reach approximately \$27.3 billion by 2030 (Grand View Research, 2023). These conversational AI agents include companionship-oriented agents (e.g., Replika (Replika, [n. d.]), Character.AI (CharacterAI, [n. d.])), mental health therapy agents (e.g, Woebot Health (Health, [n. d.]), Elomia Health (Elomia Health, [n. d.])) and general-purpose agents (e.g., OpenAI ChatGPT (Achiam et al., 2023), Microsoft Copilot (Copilot, [n. d.]), Google Gemini (Gemini, [n. d.])), with many of these already being used by individuals for specific use cases, such as supporting their well-being (Gilbert, 2024; Tidy, [n. d.]).
@@
While there is an increasing interest in the using AI agents for various tasks, previous studies have also shown that their use can pose increased risks to individuals. For instance, prior literature has examined specific risks associated AI agents, such as attachment with AI (Guerreiro and Loureiro, 2023; Pentina et al., 2023), risks from anthropomorphism (Akbulut et al., 2024), AI generating misinformative/biased text (Gravel et al., 2023; Zhou et al., 2023; Tao et al., 2023; Van Niekerk et al., 2024), or AI producing toxic content (Deshpande et al., 2023; Ovalle et al., 2023). In addition, there exists literature mapping out the space of sociotechnical risks of using generative AI into taxonomies (Shelby et al., 2023; Weidinger et al., 2022; Gabriel et al., 2024). Existing taxonomies are often developed by leveraging prior literature that identifies specific risks associated with AI (Weidinger et al., 2022; Gabriel et al., 2024) or by aligning with established theories and guidelines (Shelby et al., 2023) (e.g., feminist standpoint theory (Harding, 1991)). This approach limits their ability to uncover newer and more nuanced risks as AI agents continue to evolve and expand in their capabilities. Additionally, these taxonomies are usually not designed with the aim of building design recommendations to address the identified risks, which further limits their applicability. Additionally, prior taxonomies often present psychological risks as a sub-category within broader AI-related risks with insufficient details, leading to under-representation of the impact of psychological risks of AI use. Due to the contextual complexities involved that make psychological risks rather idiosyncratic, there is a gap in the comprehensive understanding that accurately captures, defines, and quantifies these impacts.
@@
Our work aims to fill this gap in existing AI risk taxonomies with a specific focus on psychological risks of using AI conversational agents that is grounded in lived experiences. Specifically, we take a mental-health-first perspective, gathering insights from people with lived experience in mental health as one type of “extreme user group” (Choe et al., 2014) who can provide expertise on psychological risks. Additionally, we focus on understanding the contextual complexities surrounding psychological risks, as context can influence the relationship between AI behaviors and their impact on individuals. This includes describing how behaviors lead to impact or how the context exacerbates the impact. For our work, we employed a mixed-method approach, beginning with a survey of 283 individuals with lived mental health experience, followed by three vignette-based design workshops with seven lived experience experts belonging to a longstanding technology co-design advisory board.
@@
Through our work, we make the following contributions: First, we introduce a novel psychological risk taxonomy based on the lived experiences of individuals. Our taxonomy highlights three components – (1) AI behaviors, (2) negative psychological impacts, and (3) the contexts that either mediate or exacerbate these negative impacts. Second, we propose a novel multi-path vignette framework depicting various AI behaviors, impact categories and contexts, surfacing the interplay between the three components of the taxonomy. Using this framework we gather design recommendations from the workshop participants. Lastly, using the insights from the workshop sessions, we provide practical design recommendations for AI practitioners and developers to design and develop safer AI conversational agents.
Our work holds significant importance for various stakeholders involved in examining and addressing psychological risks posed by AI agents. Our taxonomy offers valuable insights for policymakers, enabling them to create policies that prioritize a mental health-first approach. Additionally, the design recommendations provided in this work can provide actionable guidance for researchers, developers, and practitioners to create safer and more robust AI conversational agents for the future.
@@ 2. Related Work
In informing our work on building a risk taxonomy focusing on psychological risks of using AI conversational agents, we build on past research on understanding the risks posed by generative AI models, the negative psychological impacts on individuals due to interaction with technology, and studies focusing on operationalization of taxonomies for AI system design.
@@ 2.1. Risks posed by generative AI agents
With the exponential rise of generative AI based conversational agents, previous works have also studied various types of risks associated with interactions involving these systems. Past works have focused on specific harms (such as a harm to identity or information hazards) through theory-driven and algorithmic efforts stressing on quantitative evaluation. A major topic of exploration for past research has been towards understanding the impact of interactions with AI on individual well-being (De Freitas et al., 2024; Song et al., 2024; De Freitas and Cohen, 2024) and human-AI interaction specifically focusing on attachment with AI (Guerreiro and Loureiro, 2023; Pentina et al., 2023), over-reliance on AI (Jacobs et al., 2021; Chiesurin et al., 2023) and user-trust on AI (Glikson and Woolley, 2020; Chiesurin et al., 2023; Shah et al., 2022). Representational harms caused due to interaction with AI has also been studied from different facets exploring bias towards gender (Van Niekerk et al., 2024; Nadeem et al., 2021), culture (Naous et al., 2023; Tao et al., 2023), and race of individuals (Bianchi et al., 2023; Luccioni et al., 2024). Another line of research has explored risks associated with the quality and delivery of the content generated by AI conversational agents. Understanding the reliability of information has been a prominent theme among the past works particularly evaluating the presence of misinformation (Zhou et al., 2023; Achiam et al., 2023; Seow et al., 2022), inconsistency (Ohmer et al., 2023; Goyal et al., 2022) and toxicity (Deshpande et al., 2023; Ovalle et al., 2023) in content generated by AI conversational agents. Lastly, previous works have also examined the risks associated with specific AI behaviors such as sycophancy (Turpin et al., 2024; Ranaldi and Pucci, 2023; Zhao et al., 2024), manipulation (Carroll et al., 2023; Singh, 2022; Park et al., 2024), and persuasion (Dehnert and Mongeau, 2022; Kim and Duhachek, 2020). Although the exploration of specific types of harms in previous works has provided insights into the wide range of the problem, such compartmentalized view often fail to capture the plurality and diversity of experiences especially when multiple AI behaviors and impacts are related to each other in complex ways in real-world scenarios. Moreover, past studies have overlooked the contextual factors attached to users interacting with AI that can influence the severity and temporality of AI behaviors and their negative implications, further limiting their applicability to real-world scenarios.
Several works have also presented risk taxonomies that capture a broader range of impacts on individuals and society. In a recent work, researchers presented the ethical and societal risks posed by advanced AI assistants focusing on the impact of AI conversational agents over individual well-being, safety and trust (Gabriel et al., 2024). Similarly, Shelby et al. (2023) presented a taxonomy of sociotechnical harms of algorithmic harms divided into five broad categories– representational harms, allocative harms, quality-of-service harms, interpersonal harms, and social system/societal harms. In another notable past work, researchers presented categorized the risks from AI in six areas - 1) discrimination, hate speech and exclusion, 2) information Hazards, 3) misinformation harms, 4) malicious Uses, 5) human-computer interaction harms, and 6) environmental and socioeconomic harms (Weidinger et al., 2022). Additionally, past studies have also focused on addressing and evaluating a broad range of risks in AI conversational agents. In one of the recent works, researchers presented a framework for evaluating AI systems across seven categories of social impact in five modalities – text, audio, images, language and a combination of aforementioned modalities along with presenting mitigation strategies. Related to this, another popular work evaluated the trustworthiness of LLMs along seven major categories – reliability, safety, fairness, resistance to misuse, explainability and reasoning, adherence to social norms, and robustness (Liu et al., 2023b). Finally, in one of the recent works, Zhou et al. (2024) presented a risk taxonomy for LLMs in public health encompassing four dimensions – individual behaviors, human-centered care, information ecosystem, and technology accountability.
While the previous risk taxonomies provide an overview at various harmful AI behaviors and impacts, these works have mostly relied existing research which in-turn are focused on theory-driven and well-documented risks that often translates to the lack of presence of lived experiences in the the proposed taxonomies. Additionally, psychological risks have often been presented as a sub-category within broader AI-related risks in past taxonomy works, leading to neglect of under-representation of the impact of psychological risks of AI use. In contrast, our work presents a novel risk taxonomy focusing on psychological risks of using AI gathered through lived experience of individuals. Additionally, we propose a novel multi-path vignette framework for understanding the interplay between AI behaviors, negative psychological impacts and user contexts.
@@ 2.2. Past works investigating psychological impacts linked with the use of technology
Understanding and analyzing the psychological impact associated with use of technology is a widely explored area of research. In one of the early works, researchers studied technophobia among university students specifically focusing on their attitude towards the impact of technology and anxiety related to using it  (Weil and Rosen, 1995). In another early work, Billieux et al. (2007) conducted a survey among 108 psychology students evaluating the relationship between the use of and perceived dependence on the mobile phones. The authors found that the sense of urgency and lack of perseverance were linked to increased dependency on phones. Understanding the negative impacts of internet and social media platform usage has also been a central theme of past research examining the interaction between humans and technology. Prior works have explored the negative psychological impact of using internet, specifically exploring the relationship between internet addiction and psychological symptoms, such as somatization, obsessive-compulsive and depression (Adalıer and Balkan, 2012; Lin et al., 2011). Recent past works have studied various aspects related to social media use, with a particular focus on examining the negative impacts it can have on individuals. Feinstein et al. (2013) conducted a study with 268 young adults examining the impact of social comparison and found that social comparison had a significant mediating impact towards developing depressive symptoms. Other studies have explored other facets such as vaguebooking leading to suicide ideation (Berryman et al., 2018), excessive and problematic usage of social media linked with increased risk of developing loneliness (Dhir et al., 2018; Bonsaksen et al., 2023; Wu et al., 2024), depression (Reer et al., 2019; Lin et al., 2016), and social isolation (Whaite et al., 2018; Shensa et al., 2016).
Findings from past research in technology and social media literature have highlighted how the use of technology could potentially result in negative psychological implications for individuals using it. However, previous studies have also discussed how human-AI interactions differ from interpersonal human interactions due to the fundamentally distinct cognitive capabilities of AI conversational agents (Shneiderman, 2020; Korteling et al., 2021). Additionally, AI conversational agents are significantly different from past technologies due to their ability to generate content and advanced capabilities such as decomposing the tasks (Wei et al., 2022; Yao et al., 2023), advanced reasoning (Shi et al., 2022), memory of past interactions (Kim et al., 2023; Shinn et al., 2024; Liu et al., 2023a). While insights from past research on the psychological impact from the use of technology can inform our understanding, it is important to examine these implications within the context of interactions between humans and AI agents. This requires adopting a psychological risk-first perspective, focusing on the unique ways these interactions may affect human well-being.
@@ 2.3. Operationalization of taxonomies for AI design
One of the challenges with existing taxonomies is the difficulty in operationalizing them, which involves translating theoretical insights into concrete design solutions. Existing AI design pipelines focus on evaluating models using established evaluation benchmarks and frameworks. Some of the most widely adopted evaluation benchmarks used for this purpose are Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), Big-Bench-Hard (Suzgun et al., 2022), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), and AI2 Reasoning Challenge (ARC) (Clark et al., 2018). These benchmarks have been been commonly used for evaluating performance of AI conversational agents across various dimensions, such as general knowledge, reasoning, mathematical problem-solving, and commonsense understanding. While benchmarks provide an objective measure to evaluate model performance across various dimensions, they often fail to capture the nuances attached to the lived experiences of individuals, which are inherently diverse and highly context-dependent. As a result, higher scores or better performance on these benchmarks do not necessarily translate to models that are effective or equitable, as they may fail to address the specific preferences of individuals from different backgrounds and contexts.
There has been an increasing trend of conducting human evaluation on content generated by large language models to align it with individual preferences and criteria. Past works have used human evaluation for high-risk domains such as healthcare (Singhal et al., 2023; Schmidt et al., 2024; Rogasch et al., 2023), law (Savelka et al., 2023; Guha et al., 2024) and tasks such as multi-step planning (Chowdhery et al., 2023), and content safety (Ji et al., 2024). However, such evaluations are often task-specific and conducted by domain experts, excluding other stakeholders such as end-users. In contrast, our work goes beyond benchmarks and metrics by incorporating the insights and lived experiences of end-users, ensuring a more inclusive and comprehensive framework that aligns with their needs and expectations.

### v3

@@ Abstract.
Recent gains in popularity of AI conversational agents have led to their increased use for improving productivity and supporting well-being. While previous research has aimed to understand the risks associated with interactions with AI conversational agents, these studies often fall short in capturing the lived experiences of individuals. Additionally, psychological risks have often been presented as a sub-category within broader AI-related risks in past taxonomy works, leading to under-representation of the impact of psychological risks of AI use. To address these challenges, our work presents a novel risk taxonomy focusing on psychological risks of using AI gathered through the lived experiences of individuals. We employed a mixed-method approach, involving a comprehensive survey with 283 people with lived mental health experience and workshops involving experts with lived experience to develop a psychological risk taxonomy. Our taxonomy features 19 AI behaviors, 21 negative psychological impacts, and 15 contexts related to individuals. Additionally, we propose a novel multi-path vignette-based framework for understanding the complex interplay between AI behaviors, psychological impacts, and individual user contexts. Finally, based on the feedback obtained from the workshop sessions, we present design recommendations for developing safer and more robust AI agents. Our work offers an in-depth understanding of the psychological risks associated with AI conversational agents and provides actionable recommendations for policymakers, researchers, and developers.
Content Warning: This paper includes discussions of sensitive topics, including but not limited to self-harm, body shaming, and discrimination.
AI, Psychological risks, Psychological risk taxonomy, Lived experience
journalyear: 2025copyright: ccconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732063isbn: 979-8-4007-1482-5/2025/06ccs: Social and professional topics Computing / technology policyccs: Human-centered computingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Figure 1. Overview of our two-phase study and our risk taxonomy with corresponding sections in this paper. In the first phase (Section 3), we conducted a survey study (N=283), which informed the creation of the Psychological Risk Taxonomy, comprising of AI Behavior, Negative Psychological Impact, User Context, and their interplay. In the second phase (Section 4), we designed multi-path vignettes from the taxonomy and survey and conducted workshops (N=7, Sessions=3) to develop the future AI design recommendations.
[…]
This figure presents the overview of our two-phase study. In the phase 1 of the study (Section 3), we conducted a survey study (N=283), which informed the creation of the Psychological Risk Taxonomy, comprising of AI Behavior, Negative Psychological Impact, User Context, and their interplay. AI Behavior includes production of harmful content, manipulation tactics, violation of trust, and inappropriate content delivery. Negative Psychological Impact covers issues such as impact on human-AI interaction, impact on user behavior, triggering of negative emotions, harm to identity, harm to psychological safety, and mental health impacts. User Context encompasses individual, psychological, and context of use factors. We also analyzed the relationships among AI Behavior, Negative Psychological Impact, and User Context using vignettes. In the second phase (Section 4), we designed multi-path vignettes from the taxonomy and survey and conducted workshops (N=7, Sessions=3) to develop the future AI design recommendations.
@@
Since the late $\mathrm{20^{th}}$ century, advancements in technology, including personal computers and social media platforms, have enabled individuals to enhance productivity, support their well-being, and connect with other individuals (Dedrick et al., 2013; Sims et al., 2017; Hall et al., 2021; Shensa et al., 2016; Naslund et al., 2016). More recently, generative AI tools such as ChatGPT have gained popularity, which has led to an unprecedented growth in the number of individuals using these platforms to support their various productivity and well-being needs (Lebow, 2023; Afshar, 2024). Due to the natural conversational interface built on top of the underlying generative AI models that resemble human conversations, recent reports indicate that the market for conversational AI agents is projected to reach approximately \$27.3 billion by 2030 (Grand View Research, 2023). These AI conversational agents include companionship-oriented agents (e.g., Replika (Replika, [n. d.]), Character.AI (CharacterAI, [n. d.])), mental health therapy agents (e.g., Woebot Health (Health, [n. d.]), Elomia Health (Elomia Health, [n. d.])), and general-purpose agents (e.g., OpenAI ChatGPT (Achiam et al., 2023), Microsoft Copilot (Copilot, [n. d.]), Google Gemini (Gemini, [n. d.])), with many of these already being used by individuals for specific use cases, such as supporting their well-being (Gilbert, 2024; Tidy, [n. d.]).
@@
While there is an increasing interest in using AI agents for various tasks, previous studies have also shown that their use can pose increased risks to individuals, such as attachment (Guerreiro and Loureiro, 2023; Pentina et al., 2023), risks from anthropomorphism (Akbulut et al., 2024), AI generating misinformative/biased text (Gravel et al., 2023; Zhou et al., 2023; Tao et al., 2023; Van Niekerk et al., 2024), or AI producing toxic content (Deshpande et al., 2023; Ovalle et al., 2023). There also exists literature mapping out the space of sociotechnical risks of using generative AI into taxonomies (Shelby et al., 2023; Weidinger et al., 2022; Gabriel et al., 2024). These are often developed by leveraging prior literature that identifies specific risks associated with AI (Weidinger et al., 2022; Gabriel et al., 2024) or by aligning with established theories and guidelines (Shelby et al., 2023) (e.g., feminist standpoint theory (Harding, 1991)). However, such approaches have limited ability in uncovering newer and more nuanced risks as AI agents, their capabilities, and their use cases continue to evolve. These taxonomies are not designed to build AI design recommendations to address identified risks, reducing their applicability. Further, psychological risks are frequently underrepresented as they are treated as a sub-category within broader AI-related risks. The contextual complexities make psychological risks idiosyncratic, creating a gap in understanding, defining, and quantifying these impacts.
@@
Our work aims to fill this gap in existing AI risk taxonomies with a focus on psychological risks of using AI conversational agents that is grounded in lived experiences. We take a mental-health-first perspective, gathering insights from people with lived experience in mental health as one type of “extreme user group” (Choe et al., 2014) who can provide insights on psychological risks. We focus on understanding the contextual complexities surrounding psychological risks, as context can influence the relationship between AI behaviors and their impact on individuals. This includes describing how behaviors lead to impact or how the context exacerbates the impact. For our work, we employed a mixed-method approach, beginning with a comprehensive survey of 283 individuals with lived mental health experience, followed by vignette-based design workshops with seven lived experience experts.
@@
Through our work, we make the following contributions: As the primary contribution, we introduce a novel psychological risk taxonomy based on the lived experiences of individuals, highlighting three components: (1) AI behaviors, (2) negative psychological impacts, and (3) user contexts. As the secondary contribution of the work, we propose a multi-path vignette framework to demonstrate how the proposed taxonomy should be applied to surface diverse ways that AI behaviors, impact categories, and contexts interact and to inspire design recommendations. Lastly, using the insights from the workshop sessions, we provide practical design recommendations for AI practitioners and developers to design and develop safer AI conversational agents. Our work holds significant importance for various stakeholders examining and addressing psychological risks posed by AI agents. Our taxonomy offers valuable insights for policymakers to create health-first policies. Additionally, the design recommendations provide actionable guidance for researchers, developers, and practitioners to create safer AI agents.
@@ 2.1. Risks posed by generative AI agents
Past works have focused on specific harms (such as harm to identity) through theory-driven and algorithmic efforts stressing quantitative evaluation. Exploration has been towards understanding the impact of interactions with AI on individual well-being (De Freitas et al., 2024; Song et al., 2024; De Freitas and Cohen, 2024) and human-AI interaction specifically focusing on attachment with AI (Guerreiro and Loureiro, 2023; Pentina et al., 2023), over-reliance on AI (Jacobs et al., 2021; Chiesurin et al., 2023) and user trust in AI (Glikson and Woolley, 2020; Chiesurin et al., 2023; Shah et al., 2022). Representational harms caused by AI conversational agents have also been studied from different facets exploring bias towards gender (Van Niekerk et al., 2024; Nadeem et al., 2021), culture (Naous et al., 2023; Tao et al., 2023), and race of individuals (Bianchi et al., 2023; Luccioni et al., 2024). Understanding the reliability of information has also been a prominent theme among the past works, particularly evaluating the presence of misinformation (Zhou et al., 2023; Achiam et al., 2023; Seow et al., 2022), inconsistency (Ohmer et al., 2023; Goyal et al., 2022) and toxicity (Deshpande et al., 2023; Ovalle et al., 2023) in content generated by AI conversational agents. Lastly, previous works have also examined the risks associated with specific AI behaviors such as sycophancy (Turpin et al., 2024; Ranaldi and Pucci, 2023; Zhao et al., 2024), manipulation (Carroll et al., 2023; Singh, 2022; Park et al., 2024), and persuasion (Dehnert and Mongeau, 2022; Kim and Duhachek, 2020). Although the exploration of specific types of harms in previous works has provided insights into the wide range of the problem, such compartmentalized views often fail to capture the plurality and diversity of experiences, especially when multiple AI behaviors and impacts are related to each other in complex ways.
Past works have also presented risk taxonomies capturing impacts on individuals and society. A recent work presented a risk taxonomy focusing on ethical and societal risks posed by advanced AI assistants (Gabriel et al., 2024). Similarly, Shelby et al. (2023) presented a taxonomy of sociotechnical harms of algorithmic harms divided into five broad categories. In another notable past work, researchers categorized the risks from AI in six areas spanning hate speech to environmental and socioeconomic harms (Weidinger et al., 2022). Finally, in one of the recent works, Zhou et al. (2024) presented a risk taxonomy for LLMs in public health with four dimensions including individual behaviors to technology accountability. While these taxonomies provide valuable insights, they often focus on well-documented risks rather than emerging concerns inspired by lived experiences. Additionally, psychological risks are typically treated as a sub-category, leading to their under-representation. Our work addresses these gaps by presenting a novel taxonomy centered on psychological risks gathered through the lived experience of individuals.
@@ 2.2. Past works investigating psychological impacts linked with the use of technology
Understanding the psychological impact associated with technology use is an explored area of research. In one of the early works, researchers studied technophobia among university students focusing on their attitude towards the impact of technology and anxiety related to using it (Weil and Rosen, 1995). In another early work, Billieux et al. (2007) conducted a survey among 108 psychology students and observed that the sense of urgency was linked to increased dependency on phones. More recent prior works have explored the impact of using the internet and social media, specifically exploring the relationship between internet addiction and psychological symptoms, such as somatization, obsessive-compulsive disorder, and depression (Adalıer and Balkan, 2012; Lin et al., 2011)Feinstein et al. (2013) found that social comparison on social media had a significant mediating impact towards developing depressive symptoms. Other studies have explored other facets such as suicide ideation (Berryman et al., 2018), excessive and problematic usage of social media linked with increased risk of developing loneliness (Dhir et al., 2018; Bonsaksen et al., 2023; Wu et al., 2024), depression (Reer et al., 2019; Lin et al., 2016), and social isolation (Whaite et al., 2018; Shensa et al., 2016). Finally, Chen et al. (2022a) presented a framework named trauma-informed computing which incorporated six principles of trauma-informed approaches to computing namely safety, trust, peer support, collaboration, enablement, and intersectionality.
While past research in technology and social media has examined the potential negative psychological impact on individuals, previous studies have also revealed the difference between human-AI interactions and interpersonal human interactions due to the distinct cognitive capabilities of AI conversational agents (Shneiderman, 2020; Korteling et al., 2021). Further, AI conversational agents differ from past technologies due to their ability to generate content and advanced capabilities (Wei et al., 2022; Yao et al., 2023; Shi et al., 2022; Kim et al., 2023). While insights from prior research on the psychological impact of technology can provide valuable insights, understanding these effects in the context of human-AI interactions requires a psychological risk-first approach, emphasizing the unique ways these interactions may influence human well-being.
@@ 2.3. Operationalization of taxonomies for AI design
One of the challenges with existing taxonomies is the difficulty in translating theoretical insights into concrete design solutions. Existing AI design pipelines focus on evaluating models using established evaluation benchmarks and frameworks. Some of the most widely adopted evaluation benchmarks used for this purpose are Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020), Big-Bench-Hard (Suzgun et al., 2022), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), and AI2 Reasoning Challenge (ARC) (Clark et al., 2018). While benchmarks provide an objective measure to evaluate model performance across various dimensions, they often fail to capture the nuances attached to the lived experiences of individuals, which are inherently diverse and highly context-dependent. As a result, higher scores or better performance on these benchmarks do not necessarily translate to models that are effective or equitable. There has been an increasing trend of conducting human evaluation on content generated by LLMs to align it with individual preferences. Past works have used human-in-the-loop evaluation for high-risk domains such as healthcare (Singhal et al., 2023; Schmidt et al., 2024; Rogasch et al., 2023; Chandra et al., 2025), law (Savelka et al., 2023; Guha et al., 2024) and tasks such as multi-step planning (Chowdhery et al., 2023), and content safety (Ji et al., 2024). However, such evaluations are often task-specific and conducted by domain experts, excluding other stakeholders such as end-users. In contrast, our work goes beyond benchmarks and metrics by incorporating the insights and lived experiences of end-users, ensuring a more inclusive and comprehensive framework that aligns with their needs and expectations.